How do I do Elimination using Multiplacation? It is a bunch of numbers and letters with signs and I don't understand it at all. Can you help me?
Answer by elima(1433)

To solve asystem of equations you can addor subtract the two equations. You do this to "Eliminate" one variable and solvethe equation for the other. When you have solved the equation for the varialbe, then you can plug the answer into one of the priginal equation and solve for the other variable.
ex.
Lets say we are given these two equations and are told to find x& y;
2x+2y=48
x-2y=12
To do this we begin by adding the two equations together, and when you do this the y's cancel each other out;
2x+2y=48
x-2y=12
---------
3x =60
Now you can solve for x,
3x=60
x=20
Now that youhave the answer for x, you can plug that answer into the priginal equation and solve for y;
2(20)+2y=48
40+2y=48
2y=48-40
2y=8
y=4
I hope you are able to understand this explanation.
=)